OpenStudy (anonymous):

The correct solution to the radical equation is: Select one: a. 3 b. 21 c. 27 d. 29

OpenStudy (anonymous):

what is the radical equation??

OpenStudy (anonymous):

\[\sqrt{x-2}=5\]

OpenStudy (anonymous):

check

OpenStudy (solomonzelman):

square both sides

OpenStudy (anonymous):

or square and get \[x-2=25\] solve for \(x\)

OpenStudy (anonymous):

c because the square root of 25 is 5

OpenStudy (anonymous):

so square both sides... then you get x + 2 = 25 then subtract 2 then you get x = 23

OpenStudy (anonymous):

sorry i meant add 2 so then you would get 27

OpenStudy (anonymous):

lol C because it is always C or at least it is 94% of the time

OpenStudy (anonymous):

really so your saying if i guess c on my whole quiz of 10 questions ill get a 94%

OpenStudy (anonymous):

probably yes at least my experience here is that a multiple choice question has answer C way more than three quarters of the time

OpenStudy (anonymous):

satellite does have a point. C is usually the correct answer for me too. Idk why it's like that but haha. Anyways, I would go with C because I got a closer answer to C than to anything else.

OpenStudy (anonymous):

27